FEL Upgrade Project Weekly Brief
November 27 – December 1, 2000


On Nov. 27th we received from one of our generous industrial collaborators a significant equipment donation that will enhance the capabilities of the FEL User Facility:  a compact, superconducting storage ring capable of generating x-rays  with a 1nm critical wavelength.   We are currently seeking support from the Commonwealth of Virginia to construct  an addition onto the FEL building to house the synchrotron.


On Nov. 21 we distributed the monthly technical and cost/performance report for the month of October.  We received from our ONR contract monitor a draft report from the project semi-annual review that was held on Oct. 23.  To quote from the draft report:  "... plans for the 10 kW Upgrade are solid and the one major technical risk area which the panel identified at the kick-off review (the optical cavity) has been well addressed".

The project team spent two extended sessions this week discussing long range plans for the FEL, paying particular attention to next steps beyond the 10 kW device.  These discussions will be summarized for further use.

We notified two companies this week that they successfully competed for task order procurements we plan to place for engineering , design and analysis work on electron beam and optical beam transport hardware.

WBS 3 (Beam Physics):

Further progress was made in the definition of the version 1.1 lattice design. As part of the long range planning sessions mentioned above, we developed a list of beam physics issues that need to be quantified by further modeling and/or measurements on the IR Demo prior to the planned stop of machine operations approximately one year from now for upgrade installation activities.

WBS 4 (Injector):

Installed a titanium cathode into the field emission test system.  Setting up the system for a vacuum bake prior to more data collection.

WBS 6 (RF):

Zone 3 - No progress this week.

Zone 4 - All filters and cables for the waveguide couplers were installed.
              A good ADC card was found and installed.
              Hipotting was completed.  All positions were good except for #1.
              The filament board will be replaced.
              The download screens are reported to be operational.  They will be checked next week.
              Doors for the HPA have been modified and are ready for installation.
              The klystron cathode and body currents are being calibrated.  When finished the Cathode Power Supply
              (CPS) and High Power Amplifiers (HPA) will be ready for testing.

The contract for the upgrade of the cathode power supplies for the injector cryounit is ready to finalize.

WBS 8 (I&C):

Efforts were put into post-run cleanup of signal and cables. The new VME boards for beam sync distribution have been stuffed, tested, and installed. Another optical BPM was placed at the end-of-line dump, this will be tested during the run on Monday. The next cut at the closed loop system architecture was worked out with the optics group. This will
consist of five OBPM monitors on the transport controlling 3 (X&Y) mirror cans. This will be tested during the Feb. FEL run.

The beam viewer contract complete and work is progressing (vendor buying parts & materials).  The 3" BPM quotes are due Monday 12/4/00. Of 8 packages sent out we hope for 3 or 4 good quotes.

WBS 9 (Transport):

o Vendors are preparing budgetary estimates for our dipole orders.
o We continued detailing the GW dipole.
o The GW Prototype’s steel is on order.
o Results from 3D models of the GW were reviewed with G. Biallas.  It was concluded that the Purcell Gap offered no improvement to the good field and would not be required to first order and contrary to our experience.  We believe this effect is because of residual field from hysteresis.  We are making the GW prototype with provisions to have both configurations.
o The effective length and K1 were determined for all our styles of magnet and fed back to Dave Douglas for inclusion into his model and into the drawing dimensions.
o  The gap in the 180° bend was changed to 3 inches from 2 inches to reflect the thinking that beam tube dimensional changes cause wake fields that are unacceptable.
QX (3.125" Quad)
o Prototype Fabrication:  the four quadrants were received from the vendor.  Inspection and assembly of the core are underway.
o  Budgetary Quote:  three "affordable" quotes have been received for machining of the cores.  We are awaiting three additional quotes from coil vendors.
o  Measurement Probe: The 50-turn litz wire coil has been completed by the HPEE Group and fabrication of the 100-turn has started.  Drawings for the probe body are in the final steps of detailing.
o  The modifications of the Differential Pumping stations to accommodate 3 inch tube were determined
o  The shielded and unshielded bellows designs were approved.

WBS 10 (Wiggler):

The dispersion section vendor has all the materials for the magnet and has started fabrication.  The drawings for the wiggler viewers are ready for signoff.  A mistake was found in the optics calculations for the dispersion section viewer.  A new solution was found which allows us to use all the same parts for the dispersion section viewer and the
wiggler viewers.   This reduces the number of drawings necessary.  We are still waiting for the magnet group technician to be freed up to finish the wiggler assembly.  No progress this week.

WBS 11 (Optics):

Drawings for the internal components of the mirror test stand are nearly completed and will be reviewed next week.  The drawings incorporate the latest improvements on the deformable mirror we are designing under a contract with AES staff.  We also amended the statement of work to include some finite element analyses on the new design.

We installed O-BPM at the end-of-line (EOL).  The quad pyro detector electronics will be added today as well, and commissioned next week.

The custom-built enclosure for the OCR optical table was installed last week.

Changes to some of the FEL diagnostics software were implemented and will be tested next week.  These changes should make data collection easier for operators.

I prepared a presentation that was delivered at the FEL planning meetings held this week.  The presentation characterized the sort of laser parameters our users were asking for, and how these requests have evolved over the last year.


No operations for the FEL during the last two weeks.

Privacy and Security Notice